OpenStudy (idkwut):

Which of the following is a right-hand endpoint Riemann sum approximation to integral f(x)dx from 1 to 4 with six equal subdivisions?

OpenStudy (idkwut):

a. f(1)+f(1.5)+f(2)+f(2.5)+f(3)+f(3.5) b. f(1.5)+f(2)+(2.5)+f(3)+f(3.5)+f(4) c. 1/4[f(1)+f(1.5)+f(2)+f(2.5)+f(3)+(3.5)] d. 1/2[f(1)+f(1.5)+f(2)+f(2.5)+f(3)+f(3.5) e. 1/2 f(1.5)+1/2f(2) + 1/2f(2.5) + 1/2f(3) + 1/2f(3.5) + 1/2f(4)

OpenStudy (idkwut):

@myininaya

zepdrix (zepdrix):

So our interval is, \[\Large\rm (1,4)\]So let's find the width of each rectangle ( delta x ). We'll split the interval into 6 equal pieces.\[\Large\rm \Delta x=\frac{4-1}{6}\]

zepdrix (zepdrix):

So that's the width of each rectangle. 1/2 right?

OpenStudy (idkwut):

4-1/6 = 1/2, so yea

OpenStudy (idkwut):

Would it be d?

zepdrix (zepdrix):

Yes very good! Since we're using `right endpoints`, the height of our `first` rectangle is f(1.5)